EV006

PI: Volvatch
Title: OH 471

Session: Feb 1997 (session 1)
Frequency: 18 cm
Mode: B
Obs time (UT): EU 23:00(28/02)-11:00(01/03)
Stations: EbJbMcNtOnSmShUr
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (dgraham at Mon Mar 3 03:51:27 EST 1997) Failure
The operators forgot to observe this because they had an outdated list of observations.

Medicina (mauro at Mon Mar 3 22:22:55 EST 1997) Success
Clear sky and low humidity. Very good

Onsala (Antonis Polatidis at Tue Mar 4 02:44:42 EST 1997) Minor failures
Started observatins at UT 2052 when the wind came below the antenna stowing limit. Weather clear all night. No known problems with recorded data.

Effelsberg (dgraham at Tue Mar 4 03:03:35 EST 1997) Success
--last Effelsberg report wrong, just noticed it had moved by a day, took part successfully, I hope---

Shanghai (Huang Xinyong at Wed Mar 5 05:31:11 EST 1997) Success
Clock offset = 0.69 msec. No problems, weather fine.

Jodrell (Alastair Gunn at Tue Mar 18 12:54:22 MET 1997) Minor failures
Very bad interference. Vacuum problem caused loss of data after tape change until 01:26. No other known problems.

Calibration (Fredrik T Rantakyro at Thu May 8 11:38:02 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ev6.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

Medicina (Volvatch at Tue Oct 27 15:12:24 MET 1998) Minor failures